# Stability of Ecological SystemsCan Chen, Xu-Wen Wang, and Yang-Yu Liu

### 12 Dec 2023 v1

@@ 9.2.2 Tensor decomposition approach
Chen exploited tensor algebra to determine the stability of homogeneous polynomial dynamical systems [62]. First, the author proved that every $S$-dimensional homogeneous polynomial dynamical system of degree $p$ can be equivalently represented by a $(p+1)$th-order $S$-dimensional tensor, denoted by $\textsf{A}\in\mathbb{R}^{S\times S\times\stackrel{{\scriptstyle p+1}}{{\cdots}}\times S}$. Suppose that A is orthogonally decomposable, i.e.,
@@
where $\lambda_{i}$ are often referred to as the Z-eigenvalues of A with the corresponding Z-eigenvectors $\textbf{v}_{i}$ [212]. The stability properties of the homogeneous polynomial dynamical system can be readily obtained through the Z-eigenvalues, similar to linear stability. Let the initial condition $\textbf{x}_{0}=\sum_{i=1}^{S}\alpha_{i}\textbf{v}_{i}$. The equilibrium point $\textbf{x}^{*}=\textbf{0}$ is (i) stable if and only if $\lambda_{i}\alpha_{i}^{p-1}\leq 0$ for all $i$; (ii) asymptotically stable if and only if $\lambda_{i}\alpha_{i}^{p-1}<0$ for all $i$; (iii) unstable if and only if $\lambda_{i}\alpha_{i}^{p-1}>0$ for some $i$. When $p$ is odd, the initial condition can be ignored, and the stability conditions only depends on the Z-eigenvalues (exactly same as the case of linear stability).
The tensor decomposition approach provides straightforward criteria for determining the stability of homogeneous polynomial dynamical systems compared to the Lyapunov approach. Notably, Chen successfully employed this approach to analytically determine the stability of an ecological system described by the GLV model with higher-order interactions (considering third-order interactions only) with supply rates [62]. The model is defined as
@@
where $s_{i}$ is the supply rate for species $i$. The stability of the model can be obtained for orthogonal decomposable B. However, not all tensors can be decomposed in the form of (34), which limits the applicability of this approach [212]. Nevertheless, we believe that both approaches (the Lyapunov and tensor decomposition approaches) have a significant potential for analytically understanding the stability of higher-order dynamics in ecological systems.
@@ 9.3 Implicit higher-order interactions
Implicit higher-order interactions have been considered in consumer-resource models [149]. MacArthur’s consumer-resource model is one of the most commonly used models for studying consumer-resource interactions [1, 70]. It is defined as
@@
In MacArthur’s original formulation where the consumers and resources have different timescales, it can be shown that a feasible equilibrium point implies global stability by using a minimization principle [18, 25, 26]. Later on, the global asymptotic stability of MacArthur’s consumer-resource model with consumers and resources evolved in the same timescale has been proved by Case and Casten [59]. Notably, they rewrote MacArthur’s consumer-resource model into the GLV form (10) with
$\tilde{\textbf{x}}=\begin{bmatrix}x_{1}\\ \vdots\\ x_{S}\\ y_{1}\\ \vdots\\ y_{N}\end{bmatrix},\text{ }\tilde{\textbf{r}}=\begin{bmatrix}-\mu_{1}\\ \vdots\\ -\mu_{S}\\ \rho_{1}\\ \vdots\\ \rho_{N}\end{bmatrix},\text{ }\tilde{\textbf{A}}=\begin{bmatrix}0&\cdots&0&\theta_{1}\textbf{C}_{11}&\cdots&\theta_{1}\textbf{C}_{N1}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&0&\theta_{S}\textbf{C}_{1S}&\cdots&\theta_{S}\textbf{C}_{NS}\\ -\textbf{C}_{11}&\cdots&-\textbf{C}_{1S}&-\rho_{1}/K_{1}&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ -\textbf{C}_{N1}&\cdots&-\textbf{C}_{NS}&0&\cdots&-\rho_{N}/K_{N}\end{bmatrix},$
@@
Aparicio et al. explored the feasibility domain of MacArthur’s consumer-resource model [18]. Based on Case and Casten’s result, the feasibility domain can be automatically treated as the stability domain in MacArthur’s consumer-resource model. Specifically, the authors utilized the structural stability metrics as defined in (27) to compute the feasibility domain. Their findings unveil that the feasibility of MacArthur’s consumer-resource model diminishes with the pool size of the consumers increases, but the expected fraction of feasible consumers rises in this scenario [18]. Conversely, if resources exhibit linear growth, an increase in the resource pool reduces the feasibility of the model and the expected fraction of feasible consumers. However, if the resources increase logistically, the trend is reversed.
@@ 9.3.1 Impacts of cascade structure
Cascade structure plays a critical role in the stability of ecological systems [228]. Schreiber investigated the global stability of consumer-resource cascades (serially arranged containers with a dynamic consumer population that receives a flow of resources from the preceding container) [227]. The cascade model is defined as (assuming the numbers of consumers and resources are the same)
@@
for $i=1,2,\dots,S$, where $\theta$ is the assimilation efficiency of the consumer dependent on $x_{i}$, $\mu$ is the mortality rate of the consumer dependent on $x_{i}$, $\gamma$ is the flow rate, and $h(y_{i}/x_{i}^{b})$ is the functional response for $b\in[0,1]$ [23]. To ensure both biological realism and mathematical tractability, it is assumed that (i) $h:\mathbb{R}\rightarrow\mathbb{R}$ is $C^{1}$ with $h(0)=0$ and $h^{\prime}(x)>0$ for all $x\geq 0$; (ii) The limit $h_{\infty}=\lim_{x\rightarrow\infty}h(x)$ exists and is finite; (iii) $\lim_{x\rightarrow\infty}xh^{\prime}(x)=0$; (iv) $0<\theta\leq 1$ and $\mu>0$ for all $x_{i}\geq 0$; (v) $\theta^{\prime}(x_{i})\leq 0$ and $\mu^{\prime}(x_{i})\geq 0$ for all $x_{i}$; (vi) $\theta(0)h_{\infty}-\mu(0)>0$. Under these assumptions (i)-(vi), the author proved that there exists an equilibrium point of the model which is globally stable. Hence, the ratio-dependent functional response can lead to persistence of the consumer population in all containers [227].
@@ 9.3.2 Impacts of mutualistic interactions
Microbial species capable of both resource consumption and production have the potential for mutualistic interactions [75, 278]. Butler and O’Dwyer explored the stability of a consumer-resource model with mutualistic interactions by explicitly considering the resources that microbes both consume and produce [58]. Their model is defined as (assuming the numbers of consumers and resources are the same)
@@
for all $i=1,2,\dots,S$, see Fig. 14. Consequently, if microbes consume resources without producing them, any feasible equilibrium point will be locally stable. Yet, in the presence of cross-feeding, stability is no longer assured, where it can be obtained only when mutualistic interactions are sufficiently weak or when all pairs of taxa reciprocate each other’s assistance.
Figure 14: Eigenvalue distributions of the Jacobian matrix of the consumer-resource model with mutualistic interactions. (a) The randomly sampled production matrix P does not satisfy the condition (39), so there are positive eigenvalues. (b) The randomly sampled production matrix P satisfies the condition (39), so all eigenvalues are negative. This figure was drawn from [58].
@@ 9.3.3 Impacts of warming temperature
Temperature can affect the stability of ecological systems by causing changes in metabolic demands, ingestion rates, growth rates, and others [15, 16, 82, 102, 188]. Synodinos et al. used the classical Rosenzweig-MacArthur model [215] with a Holling type II functional response to study how temperature will alter stability [248]. The model is defined as (assuming one consumer and one resource)
@@ 9.3.4 Impacts of dynamic switching
The switching of the distribution of the rates of consumption between two resources is common in nature [143, 144, 145]. Gawronski et al. introduced a dynamic switching mechanism in the classical Rosenzweig-MacArthur consumer-resource model with one consumer and two resources [98]. The model is formulated as
@@ 9.3.5 Impacts of periodic environments
Periodic fluctuations and signals are widespread and strongly influence the structure and function of ecological systems [47, 217]. Bieg et al. investigated the role of periodic environmental forcing on consumer-resource interactions [51]. They used an extension of the classic Rosenzweig–MacArthur consumer-resource model [215] defined as (assuming one consumer and one resource)
@@ 9.3.6 Impacts of consumption threshold
The idea of consumption threshold is borrowed from the basic reproduction in epidemiological models [257]. Duffy and Collins introduced this notion to determine the stability of the classical Rosenzweig-MacArthur consumer-resource model [84], which is defined similarly as (assuming one consumer and one resource)
@@
The consumption threshold is defined as $C_{0}=\frac{\theta oK}{\mu(\alpha+K)}$. Ecologically, it represents the parameter combination that guarantees the minimum resource consumption required for the consumer to survive. The authors proved that the trivial equilibrium point (i.e., $x^{*}=y^{*}=0$) is always unstable, and the semi-trivial equilibrium point (i.e., $x^{*}=0$ and $y^{*}=K$) is locally stable if $C_{0}\leq 1$. Furthermore, They also demonstrated that the model undergoes a Hopf bifurcation at
@@ 9.3.7 Impacts of delayed age structure
The introduction of delay in the consumer-resource models plays a significant role in their stability. Akimenko studied the stability of a delayed age-structural consumer-resource model with $N$ resource patches and consumers that can move freely between the patches [4]. The resource dynamics of the model is defined as
@@
where $\mu(a,w)$ is the age- and calorie intake rate-dependent consumer mortality rate. Combined with initial and boundary conditions, the author proved that the trivial equilibrium point (i.e., $y_{i}^{*}=0$ and $x^{*}(a)=0$) of the model is unstable for all $\tau>0$ and the semi-trivial equilibrium point (i.e., $y_{i}^{*}=K_{i}$ and $x^{*}(a)=0$) is locally asymptotically stable if the consumer’s basic reproduction number $R(w^{*})<1$ and is unstable if $R(w^{*})\geq 1$ for all $\tau>0$. The consumer’s basic reproduction number dependent on the calorie intake rate $w$ is defined as
@@
where $\Omega(a,w)$ is the age- and calorie intake rate-dependent consumer fertility rate, and $a_{r}$ and $a_{m}$ are the age of maturation and maximum age of reproduction, respectively. Furthermore, the non-trivial equilibrium point (i.e., $y_{i}^{*}>0$ for some $i$ and $x^{*}(a)>0$) exists if and only if $R(w^{*})=1$. Define $I_{0}=\{i\text{ }|\text{ }y^{*}_{i}=0\}$. The non-trivial equilibrium point is unstable for all $\tau>0$ if $I_{0}\neq\emptyset$ and at least one of the following conditions is satisfied: (i) there exists $i\in I_{0}$ such that $\alpha_{i}\geq\beta_{i}/\rho_{i}$; or (ii) $\alpha_{i}<\beta_{i}/\rho_{i}$ for all $i\in I_{0}$ and $\hat{X}^{*}\leq\max_{i\in I_{0}}(\beta_{i}/\rho_{i}-\alpha_{i})^{-1}$. It is locally asymptotically stable for all $\tau>0$ if one of the following conditions is satisfied: (iii) $I_{0}\neq\emptyset$, $\alpha_{i}<\beta_{i}/\rho_{i}$ for all $i\in I_{0}$, and $\hat{X}^{*}>\max_{i\in I_{0}}(\beta_{i}/\rho_{i}-\alpha_{i})^{-1}$; (iv) $I_{0}=\emptyset$. Akimenko’s findings show that the digestion period (i.e., the delay parameter) $\tau$ does not destabilize the consumer population at the trivial, semi-trivial or non-trivial equilibrium points.
Recently, El-Doma improved the stability criteria proposed by Akimenko in analyzing the delayed age-structural consumer-resource model (44) and (45) [87]. Notably, the author proved that the non-trivial equilibrium point (with $I_{0}\neq\emptyset$) of the model is locally asymptotically stable if $X^{*}(\beta_{i}/\rho_{i}-\alpha_{i})>1$ and is unstable if $X^{*}(\beta_{i}/\rho_{i}-\alpha_{i})<1$ for $i\in I_{0}$. Clearly, the conditions are simpler than Akimenko’s.

### 07 Feb 2024 v2

@@ 10.2.2 Tensor decomposition approach
Chen exploited tensor algebra to determine the stability of homogeneous polynomial dynamical systems chen2022explicit . First, the author proved that every $S$-dimensional homogeneous polynomial dynamical system of degree $p$ can be equivalently represented by a $(p+1)$th-order $S$-dimensional tensor, denoted by $\textsf{A}\in\mathbb{R}^{S\times S\times\stackrel{{\scriptstyle p+1}}{{\cdots}}\times S}$. Suppose that A is orthogonally decomposable, i.e.,
@@
where $\lambda_{i}$ are often referred to as the Z-eigenvalues of A with the corresponding Z-eigenvectors $\textbf{v}_{i}$ robeva2016orthogonal . The stability properties of the homogeneous polynomial dynamical system can be readily obtained through the Z-eigenvalues, similar to linear stability. Let the initial condition $\textbf{x}_{0}=\sum_{i=1}^{S}\alpha_{i}\textbf{v}_{i}$. The equilibrium point $\textbf{x}^{*}=\textbf{0}$ is (i) stable if and only if $\lambda_{i}\alpha_{i}^{p-1}\leq 0$ for all $i$; (ii) asymptotically stable if and only if $\lambda_{i}\alpha_{i}^{p-1}<0$ for all $i$; (iii) unstable if and only if $\lambda_{i}\alpha_{i}^{p-1}>0$ for some $i$. When $p$ is odd, the initial condition can be ignored, and the stability conditions only depends on the Z-eigenvalues (exactly same as the case of linear stability).
The tensor decomposition approach provides straightforward criteria for determining the stability of homogeneous polynomial dynamical systems compared to the Lyapunov approach. Notably, Chen successfully employed this approach to analytically determine the stability of an ecological system described by the GLV model with higher-order interactions (considering third-order interactions only) with supply rates chen2022explicit . The model is defined as
@@
where $s_{i}$ is the supply rate for species $i$. The stability of the model can be obtained for orthogonal decomposable B. However, not all tensors can be decomposed in the form of (36), which limits the applicability of this approach robeva2016orthogonal . Nevertheless, we believe that both approaches (the Lyapunov and tensor decomposition approaches) have a significant potential for analytically understanding the stability of higher-order dynamics in ecological systems.
@@ 10.3 Implicit higher-order interactions
Implicit higher-order interactions have been considered in consumer-resource models lafferty2015general . MacArthur’s consumer-resource model is one of the most commonly used models for studying consumer-resource interactions advani2018statistical ; chesson1990macarthur . It is defined as
@@
In MacArthur’s original formulation where the consumers and resources have different timescales, it can be shown that a feasible equilibrium point implies global stability by using a minimization principle aparicio2023feasibility ; arthur1969species ; arthur1970species . Later on, the global asymptotic stability of MacArthur’s consumer-resource model with consumers and resources evolved in the same timescale has been proved by Case and Casten case1979global . Notably, they rewrote MacArthur’s consumer-resource model into the GLV form (10) with
$\footnotesize\tilde{\textbf{x}}=\begin{bmatrix}x_{1}\\ \vdots\\ x_{S}\\ y_{1}\\ \vdots\\ y_{N}\end{bmatrix},\text{ }\tilde{\textbf{r}}=\begin{bmatrix}-\mu_{1}\\ \vdots\\ -\mu_{S}\\ \rho_{1}\\ \vdots\\ \rho_{N}\end{bmatrix},\text{ }\tilde{\textbf{A}}=\begin{bmatrix}0&\cdots&0&\theta_{1}\textbf{C}_{11}&\cdots&\theta_{1}\textbf{C}_{N1}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&0&\theta_{S}\textbf{C}_{1S}&\cdots&\theta_{S}\textbf{C}_{NS}\\ -\textbf{C}_{11}&\cdots&-\textbf{C}_{1S}&-\rho_{1}/K_{1}&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ -\textbf{C}_{N1}&\cdots&-\textbf{C}_{NS}&0&\cdots&-\rho_{N}/K_{N}\end{bmatrix},$
@@
Aparicio et al. explored the feasibility domain of MacArthur’s consumer-resource model aparicio2023feasibility . Based on Case and Casten’s result, the feasibility domain can be automatically treated as the stability domain in MacArthur’s consumer-resource model. Specifically, the authors utilized the structural stability metrics as defined in (29) to compute the feasibility domain. Their findings unveil that the feasibility of MacArthur’s consumer-resource model diminishes with the pool size of the consumers increases, but the expected fraction of feasible consumers rises in this scenario aparicio2023feasibility . Conversely, if resources exhibit linear growth, an increase in the resource pool reduces the feasibility of the model and the expected fraction of feasible consumers. However, if the resources increase logistically, the trend is reversed.
@@ 10.3.1 Impacts of cascade structure
Cascade structure plays a critical role in the stability of ecological systems shanafelt2018stability . Schreiber investigated the global stability of consumer-resource cascades (serially arranged containers with a dynamic consumer population that receives a flow of resources from the preceding container) schreiber1996global . The cascade model is defined as (assuming the numbers of consumers and resources are the same)
@@
for $i=1,2,\dots,S$, where $\theta$ is the assimilation efficiency of the consumer dependent on $x_{i}$, $\mu$ is the mortality rate of the consumer dependent on $x_{i}$, $\gamma$ is the flow rate, and $h(y_{i}/x_{i}^{b})$ is the functional response for $b\in[0,1]$ arditi1991functional . To ensure both biological realism and mathematical tractability, it is assumed that (i) $h:\mathbb{R}\rightarrow\mathbb{R}$ is $C^{1}$ with $h(0)=0$ and $h^{\prime}(x)>0$ for all $x\geq 0$; (ii) The limit $h_{\infty}=\lim_{x\rightarrow\infty}h(x)$ exists and is finite; (iii) $\lim_{x\rightarrow\infty}xh^{\prime}(x)=0$; (iv) $0<\theta\leq 1$ and $\mu>0$ for all $x_{i}\geq 0$; (v) $\theta^{\prime}(x_{i})\leq 0$ and $\mu^{\prime}(x_{i})\geq 0$ for all $x_{i}$; (vi) $\theta(0)h_{\infty}-\mu(0)>0$. Under these assumptions (i)-(vi), the author proved that there exists an equilibrium point of the model which is globally stable. Hence, the ratio-dependent functional response can lead to persistence of the consumer population in all containers schreiber1996global .
@@ 10.3.2 Impacts of mutualistic interactions
Microbial species capable of both resource consumption and production have the potential for mutualistic interactions coyte2015ecology ; zelezniak2015metabolic . Butler and O’Dwyer explored the stability of a consumer-resource model with mutualistic interactions by explicitly considering the resources that microbes both consume and produce butler2018stability . Their model is defined as (assuming the numbers of consumers and resources are the same)
@@
for all $i=1,2,\dots,S$, see Figure 15. Consequently, if microbes consume resources without producing them, any feasible equilibrium point will be locally stable. Yet, in the presence of cross-feeding, stability is no longer assured, where it can be obtained only when mutualistic interactions are sufficiently weak or when all pairs of taxa reciprocate each other’s assistance.
Figure 15: Eigenvalue distributions of the Jacobian matrix of the consumer-resource model with mutualistic interactions. (a) The randomly sampled production matrix P does not satisfy the condition (41), so there are positive eigenvalues. (b) The randomly sampled production matrix P satisfies the condition (41), so all eigenvalues are negative. This figure was drawn from butler2018stability .
@@ 10.3.3 Impacts of warming temperature
Temperature can affect the stability of ecological systems by causing changes in metabolic demands, ingestion rates, growth rates, and others amarasekare2015effects ; amarasekare2019effects ; donohue2016navigating ; gillooly2001effects ; pawar2016real . Synodinos et al. used the classical Rosenzweig-MacArthur model rosenzweig1963graphical with a Holling type II functional response to study how temperature will alter stability synodinos2023effects . The model is defined as (assuming one consumer and one resource)
@@ 10.3.4 Impacts of dynamic switching
The switching of the distribution of the rates of consumption between two resources is common in nature kvrivan1996optimal ; kvrivan2003optimal ; kvrivan1999optimal . Gawronski et al. introduced a dynamic switching mechanism in the classical Rosenzweig-MacArthur consumer-resource model with one consumer and two resources gawronski2022instability . The model is formulated as
@@ 10.3.5 Impacts of periodic environments
Periodic fluctuations and signals are widespread and strongly influence the structure and function of ecological systems bernhardt2020life ; rubin2023irregular . Bieg et al. investigated the role of periodic environmental forcing on consumer-resource interactions bieg2023stability . They used an extension of the classic Rosenzweig–MacArthur consumer-resource model rosenzweig1963graphical defined as (assuming one consumer and one resource)
@@ 10.3.6 Impacts of consumption threshold
The idea of consumption threshold is borrowed from the basic reproduction in epidemiological models van2002reproduction . Duffy and Collins introduced this notion to determine the stability of the classical Rosenzweig-MacArthur consumer-resource model duffy2016identifying , which is defined similarly as (assuming one consumer and one resource)
@@
The consumption threshold is defined as $C_{0}=\frac{\theta oK}{\mu(\alpha+K)}$. Ecologically, it represents the parameter combination that guarantees the minimum resource consumption required for the consumer to survive. The authors proved that the trivial equilibrium point (i.e., $x^{*}=y^{*}=0$) is always unstable, and the semi-feasible equilibrium point (i.e., $x^{*}=0$ and $y^{*}=K$) is locally stable if $C_{0}\leq 1$. Furthermore, They also demonstrated that the model undergoes a Hopf bifurcation at
@@ 10.3.7 Impacts of delayed age structure
The introduction of delay in the consumer-resource models plays a significant role in their stability. Akimenko studied the stability of a delayed age-structural consumer-resource model with $N$ resource patches and consumers that can move freely between the patches akimenko2021stability . The resource dynamics of the model is defined as
@@
where $\mu(a,w)$ is the age- and calorie intake rate-dependent consumer mortality rate. Combined with initial and boundary conditions, the author proved that the trivial equilibrium point (i.e., $y_{i}^{*}=0$ and $x^{*}(a)=0$) of the model is unstable for all $\tau>0$ and the semi-feasible equilibrium point (i.e., $y_{i}^{*}=K_{i}$ and $x^{*}(a)=0$) is locally asymptotically stable if the consumer’s basic reproduction number $R(w^{*})<1$ and is unstable if $R(w^{*})\geq 1$ for all $\tau>0$. The consumer’s basic reproduction number dependent on the calorie intake rate $w$ is defined as
@@
where $\Omega(a,w)$ is the age- and calorie intake rate-dependent consumer fertility rate, and $a_{r}$ and $a_{m}$ are the age of maturation and maximum age of reproduction, respectively. Furthermore, the non-trivial equilibrium point (i.e., $y_{i}^{*}>0$ for some $i$ and $x^{*}(a)>0$) exists if and only if $R(w^{*})=1$. Define $I_{0}=\{i\text{ }|\text{ }y^{*}_{i}=0\}$. The non-trivial equilibrium point is unstable for all $\tau>0$ if $I_{0}\neq\emptyset$ and at least one of the following conditions is satisfied: (i) there exists $i\in I_{0}$ such that $\alpha_{i}\geq\beta_{i}/\rho_{i}$; or (ii) $\alpha_{i}<\beta_{i}/\rho_{i}$ for all $i\in I_{0}$ and $\hat{X}^{*}\leq\max_{i\in I_{0}}(\beta_{i}/\rho_{i}-\alpha_{i})^{-1}$. It is locally asymptotically stable for all $\tau>0$ if one of the following conditions is satisfied: (iii) $I_{0}\neq\emptyset$, $\alpha_{i}<\beta_{i}/\rho_{i}$ for all $i\in I_{0}$, and $\hat{X}^{*}>\max_{i\in I_{0}}(\beta_{i}/\rho_{i}-\alpha_{i})^{-1}$; (iv) $I_{0}=\emptyset$. Akimenko’s findings show that the digestion period (i.e., the delay parameter) $\tau$ does not destabilize the consumer population at the trivial, semi-feasible or non-trivial equilibrium points.
Recently, El-Doma improved the stability criteria proposed by Akimenko in analyzing the delayed age-structural consumer-resource model (46) and (47) el2023age . Notably, the author proved that the non-trivial equilibrium point (with $I_{0}\neq\emptyset$) of the model is locally asymptotically stable if $X^{*}(\beta_{i}/\rho_{i}-\alpha_{i})>1$ and is unstable if $X^{*}(\beta_{i}/\rho_{i}-\alpha_{i})<1$ for $i\in I_{0}$. Clearly, the conditions are simpler than Akimenko’s.